# ADA-SAM: Self-Prompting, Self-Correcting Segment Anything Model for the Analysis of Muscle Dysfunction

### v1

@@ Abstract
Keywords:
Medical imaging MRI Multi-task learning SAM Semi-supervised learning.
@@ 2.2 SegEx: Expert-driven Segmentation Assessment
To evaluate the clinical applicability and reliability of ADA-SAM in segmenting the VL and VM muscles, we introduce an innovative expert assessment (SegEx). SegEx is conducted by setting up a visual Turing test employing expert observers — human or large language model (LLM) observers. The goal is to assess the quality of the ADA-SAM generated masks for their clinical utility. First, consulting with domain expert collaborators, we select a set of clinically-relevant evaluation criteria $E=\{e_{1},e_{2},\dots,e_{k}\}$. As defined in Table 1, the selected criteria focus on mask quality, size difference, whether any correction is needed, and diagnostic confidence. Each of the criteria may have different scales; they are standardized ($\hat{e}_{i}$) to be in a common range $[r_{min}$, $r_{max}]$. Any binary decision (0/1) criteria are multiplied by the maximum $r_{max}$. The final evaluation score is then obtained by calculating the average across all the criteria, i.e., $E_{avg}=\frac{1}{k}\sum^{k}_{1}\hat{e}$.
@@ 3.1 Implementation Details
Data: We validated our proposed ADA-SAM method on a dataset obtained from individuals with patellar instability [1]. This is a multi-class dataset containing segmentation information for two quadriceps muscles: VL and VM. The data was collected from a Turbo Spin Echo MRI sequence in 65 slice stacks with a field of view of 192$\times$192 mm, an acquisition matrix of 256$\times$256, 6 mm slice thickness, and a voxel size of 0.75$\times$0.75$\times$6 mm. Originally, the dataset contained scans from 21 patients, which we split into train, validation, and test sets using a 15/1/5 split containing 975/65/325 slices, respectively.
@@
#Slices Class UMTL UMTL++ nnUMTL SAM-Mix SAM-PP ADA-SAM* ADA-SAM

### v2

# Autoadaptive Medical Segment Anything Model

@@ 2.2 SegEx: Expert-driven Segmentation Assessment
To evaluate the clinical applicability and reliability of ADA-SAM in segmenting the VL and VM muscles, we introduce an innovative expert assessment (SegEx). SegEx is conducted by setting up a visual Turing test employing expert observers — human or large language model (LLM) observers. The goal is to assess the quality of the ADA-SAM generated masks for their clinical utility. First, consulting with domain expert collaborators, we select a set of clinically-relevant evaluation criteria $E=\{e_{1},e_{2},\dots,e_{k}\}$. As defined in Table 1, the selected criteria focus on mask quality, size difference, whether any correction is needed, and diagnostic confidence. Each of the criteria may have different scales; they are standardized ($\hat{e}_{i}$) to be in a common range $r_{min}$, $r_{max}]$. Any binary decision (0/1) criteria are multiplied by the maximum $r_{max}$. The final evaluation score is then obtained by calculating the average across all the criteria, i.e., $E_{avg}=\frac{1}{k}\sum^{k}_{1}\hat{e}$.
@@ 3.1 Implementation Details
Data: We validated our proposed ADA-SAM method on a dataset obtained from individuals with patellar instability [1]. This is a multi-class dataset containing segmentation information for two quadriceps muscles: VL and VM. The data was collected from a Turbo Spin Echo MRI sequence in 65 slice stacks with a field of view of $\times192$ mm, an acquisition matrix of $\times256$, 6 mm slice thickness, and a voxel size of $\times0.75$$\times$6 mm. Originally, the dataset contained scans from 21 patients, which we split into train, validation, and test sets using a 15/1/5 split containing 975/65/325 slices, respectively.
@@
#Slices Class UMTL UMTL++ nnUMTL SAM-Mix SAM-$P$ ADA-SAM* ADA-SAM